OpenStudy (anonymous):

Find the y intercept in the following equation: y=3/4x+10 0,3/4 3/4,0 0,10 10,0

OpenStudy (amistre64):

what is the value of x for all points in the y axis?

OpenStudy (anonymous):

that's wat im trying to find

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@BrianMatthew

OpenStudy (texaschic101):

when you have an equation in y = mx + b form, the b position is the y intercept. y = mx + b y = 3/4x + 10 now what number is in the b position ?

OpenStudy (texaschic101):

any thoughts ?

OpenStudy (anonymous):

nope sorry

OpenStudy (texaschic101):

look at this formula : y = mx + b now look at your equation : y = 3/4x + 10 you can see that the m position is 3/4, by the way, that is the slope. Now what number is in the b position ?

OpenStudy (texaschic101):

 now what number is in the b position ?